OpenStudy (anonymous):

the function f is given by the formula ƒ(x)= 4x^3-3x^2+7x-8÷ x-1 when x<1 and by the formula ƒ(x)=-4x^2-2x+a when 1≤x What value must be chosen for a in order to make this function continuous at 1? a=______?? My answer is -1 but not sure if it's correct!

OpenStudy (mertsj):

I got 19. How did you get -1?

OpenStudy (anonymous):

For f(x) to be continuous at x=1 left hand limit of f at x=1 should be equal to the right hand limit at x=1 equal to f(1) f(1) = 4-2+a = 2+a left hand limit of f at x=1 is f(1-h) = limh→0 ( 4(1-h)^3-3(1-h)^2+7(1-h)-8)/h since h tends to zero so we can use binomial expansion (1-h)^n = 1-nh-n*n-1*h*h/2........ Neglecting higher powers... f(1-h) = (4(1-3h)-3(1-2h)+7+7h-8)/h f(1-h) = (4-12h-3+6h+7+7h-8)/h = h/h =1 f(1) = 2+a = f(1-h) = 1 2+a = 1 Thus a= -1 Nooo??